OpenStudy (anonymous):

how do I solve sqrt8^3+sqrt2x

OpenStudy (anonymous):

\[\sqrt{8^{3}}+\sqrt{2x}\]?

OpenStudy (anonymous):

That's not an equation, so it cannot be solved. It can be simplified however. Is that what you need?

OpenStudy (anonymous):

haha my bad yes

OpenStudy (anonymous):

Ok. First, what is 8^3?

OpenStudy (anonymous):

22.62

OpenStudy (anonymous):

That is not 8^3. That is 8^(3/2) truncated at two decimal places. Again, what is 8^3?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

*Sigh.* \[8^{3}=8\times8\times8\]What is that equal to?

OpenStudy (anonymous):

512

OpenStudy (anonymous):

Right. Now\[\sqrt{512}=\sqrt{2\times 256}=\sqrt{2}\times\sqrt{256}\]What is\[\sqrt{256}\]?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

Right. So\[\sqrt{8^{3}}+\sqrt{2x}=16\sqrt{2}+x^{\frac{1}{2}}\sqrt{2}=(16+x^{\frac{1}{2}})\sqrt{2}\]